The sum of two numbers is 21. Their product is 110. Find the numbers
The two numbers are____ and ____.

Answer by mananth(16946)

If one number is x the other is 21-x
x(21-x) = 110
21x-x^2=110
x^2-21x+110=0
(x-10)(x-11) =0
x = 10 and 11